A lad agreed to work one year for $240.00 and a calf. At the end of seven months he quit and received $100.00 and the calf. What was the value of the calf?
Answer by checkley71(8403)

7/12(240+C)=100+C OR 140+(7/12)C=100+C OR 140-100=C-(7/12)C OR 40=(5/12)C OR
40*12=5C OR 480=5C OR C=480/5 OR C=$96 IS THE VALUE OF THE CALF.